LASER SPRECTROSCOPIC STUDIES OF Si_{3}C AND Si_{3}

\underline{J. DUDEK}, M. C. MCCARTHY and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and Division of Engineering Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138; J. F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712.

Laser experiments to measure the electronic spectra of Si_{3}C and Si_{3} in the gas phase by R2C2PI spectroscopy with Time of Flight Mass Spectrometry will be presented. The strong C̃^{1}B_{1} ⟵ X̃^{1}A_{1} transition of Si_{3}C is detected at 24933 cm^{−1}. The complex band profile can be explained as vibrational progressions from at least two modes. Using matrix data from J. Fulara et al.^{1}, the ^{1}A_{1} ⟵ X̃^{1}A_{1} band of Si_{3} at 18605 cm^{−1} has also been observed. The status of the search for the microwave spectrum of Si_{3}C on the basis of recent high level ab initio calculations will be discussed. Both molecules are of astronomical interest because they are structurally similar to the known astronomical SiC_{2} and SiC_{3} species.

1 J. Fulara, P. Freivogel, M. Grutter, and J. P. Maier, J. Phys. Chem. 100, 18042 (1996).